Window problem (2 P) Consider a skyscraper in a hurricane. The wind reaches v -144 km/h and presses against a 1.25m x 2.5m-window Wall Window Assume the air to be incompressible with a density of p1.2kg/m3. km -144 Calculate the winds force on the window! Water Tank and block (3 P) Water flows from a large tank through an orifice of 4 cm diameter and against a block Water H tank The block weighs 10 kg, and the coefficient of friction between block and floor is 0,57 Block G-10kg What is the minimum height to which the water must rise in the tank in order to start the block moving to the right? h, 5cm
Answer #1

Problem 2P:

the velocity of wind, v = 144 km/h = (144 * (5/18)) m/s = 40 m/s

density of air, p= 1.2 kg / m3

Pressure, P = 0.5 * (p * v2)

Area of the window, A = (1.25 * 2.5) m2

Wind's force on the window, F = P*A = 0.5 * (p * v2) *A = 3000 N = 3 kN (ans)